# Grounded 3D-LLM with Referent Tokens

### v1

@@ Abstract
Prior studies on 3D scene understanding have primarily developed specialized models for specific tasks or required task-specific fine-tuning. In this study, we propose Grounded 3D-LLM, which explores the potential of 3D large multi-modal models (3D LMMs) to consolidate various 3D vision tasks within a unified generative framework. The model uses scene referent tokens as special noun phrases to reference 3D scenes, enabling the handling of sequences that interleave 3D and textual data. It offers a natural approach for translating 3D vision tasks into language formats using task-specific instruction templates. To facilitate the use of referent tokens in subsequent language modeling, we have curated large-scale grounded language datasets that offer finer scene-text correspondence at the phrase level by bootstrapping existing object labels. Subsequently, we introduced Contrastive LAnguage-Scene Pre-training (CLASP) to effectively leverage this data, thereby integrating 3D vision with language models. Our comprehensive evaluation covers open-ended tasks like dense captioning and 3D QA, alongside close-ended tasks such as object detection and language grounding. Experiments across multiple 3D benchmarks reveal the leading performance and the broad applicability of Grounded 3D-LLM. Code and datasets will be released on the project page.
@@ 1 Introduction
However, the exploration of a unified model for 3D perception and reasoning remains scarce, leaving it as an open problem. In our study, by leveraging the generative modeling capability of large language models (LLM), we introduce Grounded 3D-LLM for integration of multiple 3D vision tasks in linguistic formats (Fig. 1). To connect perception and language, this model represents scene regions or object features as a special token, termed “referent tokens” (<ref>). Referent tokens support decoding to identify their corresponding entities in the scene. For instance, <ref> can denote entities such as “the nearby nightstand” or “three brown chairs”, substituting the corresponding text tokens in the language. The use of referent tokens as manageable noun phrases in linguistic descriptions includes two main features: (i) A referent token can either replace or accompany text phrases, as depicted in Fig. 1. (ii) A referent token can be decoded into scene entities, encompassing single or multiple objects in a 3D scene, using simple decoder heads. With this design, referent tokens serve as the unified interface, facilitating 3D grounding capabilities in various 3D vision tasks.
The referent token in our language modeling relies on phrase-level scene-text correspondence, setting it apart from models like CLIP [57], which align at the sentence level. To achieve this, inspired by phrase-to-region correspondence [40], we employ Contrastive LAnguage-Scene Pre-training (CLASP) for a more granular scene-text correspondence. Specifically, CLASP aligns noun phrases with visual queries [62] representing scene objects or regions. However, large-scale paired scene-text data at the phrase level remains scarce. To address this, we developed an automated pipeline for annotating scene-level language data, utilizing densely-annotated object labels from indoor detection datasets to create the Grounded Scene Caption (G-SceneCap) data. This dataset comprises 107K local scene captions with grounding, offering high object density with an average of one instance every nine tokens as shown in Table 2. Additionally, we converted existing human-annotated language data [8, 83] into grounded language forms for training and evaluation. Our experiments using this data in supervised and zero-shot settings [88] show broad generalization across grounding benchmarks.
After the pre-training stage, we conduct multi-task instruction tuning for Grounded 3D-LLM using referent tokens. This model supports the decoding of these tokens into mask embeddings, enabling the retrieval of point-level masks and thus identifying their external scene referents. It supports a range of existing 3D vision tasks—from single to multi-object grounding, instance segmentation, to 3D QA and captioning—by employing diverse templates of task-specific instruction-tuning, as shown in Fig. 1. We compare this model’s prompt support and task versatility with previous large multi-modal models in Table 1. We believe this represents a valid step toward unifying various 3D vision tasks within generative language modeling.
@@
Figure 1: Unifying scene-level 3D vision tasks via task-specific prompts in Grounded 3D-LLM framework. The interleaved referent token <ref> in the language output enables decoding for point-level masks. Matching colors are shown in the figure to connect phrases with their respective object masks.
@@
  • We developed Grounded 3D-LLM, which first establishes a correspondence between 3D scenes and language phrases through referent tokens. This method enhances scene referencing and effectively supports 3D vision tasks in language modeling, including single- and multi-object grounding, along with introducing 3D detection for the first time.
  • We developed a carefully crafted automated 3D scene caption dataset curation pipeline that provides finer correspondence at the phrase level. Experiments using CLASP in both supervised and zero-shot text settings demonstrate the effectiveness of pre-training on this data for phrase-level scene-text alignment.
  • The Grounded 3D-LLM model tackles 3D grounding and language tasks generatively without the need for specialized models. It achieves top-tier performance in most downstream tasks among generative models, particularly in grounding problems, without task-specific fine-tuning.
@@
Text Vision Inst.Seg. Obj.Det. Grd. Point-Grd. Multi-Obj Grd. QA Cap
@@
Table 1: Comparison of 3D multi-modal models. We refer to instance segmentation as Inst.Seg., object box detection as Obj.Det., single-object grounding as Grd., point-level grounding as Point-Grd., multi-object grounding as Multi-Obj Grd., question answering as QA, and dense captioning as Cap.
@@ 2 Related Works
2D large multi-modal models. Vision-language models [57, 44, 43] are proposed to bridge visual and textual models via image-level contrastive learning. Their direct application enables open-world understanding [28, 80]. GLIP and its subsequent works [45, 82, 77, 49] have innovatively formulated object detection as grounding problems, leveraging detection or pseudo-labels to align semantics. With stronger language models, various powerful large multi-modal models are capable of handling tasks including image captioning [35, 54, 3], visual question answering [47, 87, 17, 70], and object detection/segmentation [54, 49, 58]. Besides, incorporating multi-modal tokens into language modeling has been validated, including visual understanding [3, 69, 42, 81], robot control [59, 5], and interleaved text-image content generation [63, 25, 26].
@@
3D scene understanding models. In the domain of 3D scene comprehension, language is progressively utilized as a contextual tool to encapsulate user intentions, given its extensive applicability. This spans from close-set object detection [37, 50, 62, 68, 36] to open-set 3D detection [10, 22, 23, 86]. Typically, these 3D detectors lay the groundwork for subsequent scene-language tasks. Such tasks encompass 3D Visual Grounding [8, 2, 34, 83, 12, 71, 84], where models pinpoint objects within a 3D scene based on language queries. They also include 3D Question Answering [4, 53, 51], which demands robust spatial perception and reasoning. Another task, 3D Dense Captioning [14, 79, 38], involves localizing and describing all scene objects, necessitating a comprehensive understanding of object locations and attributes. While initial efforts were often task-specific and lacked generalizability, some methodologies [6, 9] have integrated 3D visual grounding and dense captioning tasks, leveraging their complementary nature. Recent endeavors such as 3D-VisTA [88] and 3D-VLP [39] aim to establish a universal 3D visual-language framework by pre-aligning 3D scenes with language. In contrast to our model, most still concentrate on close-set scene understanding, striving to develop specialist models or requiring per-task fine-tuning. Moreover, they typically rely on prior detectors.
@@
3D large multi-modal models. 3D object-level LMMs [78, 74, 75, 48, 85] have effectively bridged the gap between 3D visuals and texts by leveraging extensive 3D object datasets [19, 18]. This has advanced the development of 3D object-level LMM [30, 29, 73, 31, 56, 55, 85, 46]. However, these models struggle with interpreting complex spatial interrelationships in 3D scenes. For scene-level LLMs, Chat-3D [72] and LL3DA [11] facilitate interaction with scene objects via pre-selection methods. Chat-3D v2 [33] enhances this by using object identifiers for precise referencing and grounding. 3D-LLM [31] improves scene understanding by incorporating positional embeddings and location tokens. Our approach utilizes special referent tokens for flexible 3D scene modeling and supports various tasks including object detection and grounding. Additionally, our model functions as a generalist model without requiring task-specific fine-tuning.
@@ 3 Method
This section is structured into three main parts. Initially, Sec. 3.1 explores different levels of vision-language alignment across sentences, words, and phrases. Next, Sec. 3.2 details the Grounded 3D-LLM, from vision-language pre-training to multi-task instruction-tuning using referent tokens. Finally, Sec. 3.3 describes construction of the grounded language data.
@@ 3.1 Preliminary: Vision-Language Correspondence
The significance of vision-language alignment in 2D foundation models, exemplified by the Contrastive Language-Image Pre-training (CLIP) model [61], is well-established. We explore three levels of vision-language alignment:
  • Sentence-level Correspondence: As exemplified by CLIP models [57, 44, 43] and 3D annotation works [8, 2, 83], this constraint aligns image-text pairs at the sentence level. However, it is limited by the scarcity of vision-language data, particularly in 3D.
  • Word-level Correspondence: This term describes the method of concatenating detection labels into a sentence, rather than embedding phrases within sentences, first introduced by GLIPs [45, 82, 77]. This approach reinterprets object detection labels as phrase grounding tasks. However, it typically lacks detailed appearance annotations and does not capture spatial relationships.
  • Phrase-level Correspondence: As referred to as phrase-to-region correspondence in prior work [40], this constraint requires linking text phrases to real-world objects within a context description. Correspondences may be one-to-one or multiple.
Our motivation. Phrase-to-region correspondence provides a more granular scene-text alignment than sentence-level correspondence, enabling effective modeling of grouping objects with similar attributes such as spatial relationships or appearances within the text context. A vision-language model is first pre-trained on extensive scene-text data to establish this phrase-level cross-modal correspondence. Utilizing this semantically enriched scene representation, an LLM model is then trained to integrate referent tokens into object clusters with shared linguistic descriptions. This process aligns the hidden embeddings of referent tokens with their corresponding visual referents, mirroring the phrase-to-region approach. For example, in phrases like “two brown chairs nearby <ref>” or simply “object <ref>”, a mask decoder can be trained to produce mask embeddings of <ref> to deduce the positions and shapes of these target chairs.
@@ 3.2 Grounded 3D-LLM
The main pipeline of Grounded 3D-LLM is depicted in Fig. 2. In the pre-training stage, contrastive language-scene pre-training (CLASP) is performed to align point and text embeddings at the phrase level on the extensive scene-text data. During the instruction-tuning phase, the overall Grounded 3D-LLM is trained to utilize referent tokens from the scene as soft text tokens to tackle various 3D vision tasks in a unified generative modeling framework.
Step 1: Contrastive Language-Scene Pre-training. Grounded 3D-LLM relies on referent tokens to reference the 3D scene, necessitating phrase-level alignment between natural phrases and scene referents. To bridge 3D and text modality gap, we propose the Contrastive Language-Scene Pre-training (CLASP). This model co-trains a 3D encoder, a text encoder, and a cross-modal interactor for text-guided instance segmentation as the left side of Fig. 2. In the CLASP’s implementation, a sparse convolutional network [27] serves as the 3D encoder to extract point-level embeddings $P$, while BERT [21] is utilized for extracting text embeddings $T$. Utilizing these embeddings, the cross-modal interactor employs $K$ learnable queries $Q$ as proxies to connect text with raw point clouds. These queries interact with point embeddings through cross-attention modules and with text embeddings via bidirectional cross-attention modules [49]. The model is initialized with a pre-trained 3D detector [62] to accelerate the convergence of queries serving as object proposals. At last, this multi-modal model outputs two text embeddings: $P_{text}$ for positive text tokens and $Q_{text}$ for scene queries, along with $P_{mask}$ and $Q_{mask}$, the mask embeddings for points and queries, respectively:
$\left[P_{mask},Q_{mask}\right],\left[Q_{text},P_{text}\right]=\text{Cross-Modal Interactor}(P,T,Q).$ (1)
@@
As depicted in Fig. 2, we compute the alignment scores $S_{mask}$ for per-point mask classification and similarity $S_{text}$ between text phrases and their corresponding entities (could be multiple instances):
$S_{mask}=P_{mask}Q_{mask}^{\top},\ \ S_{text}=P_{text}Q_{text}^{\top}.$ (2)
@@
The contrastive loss for $S_{mask}$ is calculated based on assignments using bipartite matching. Upon establishing the matching indices for the mapping between instances and queries, another contrastive loss for noun phrases and instance alignment, $S_{text}$ is supervised. Both similarity matrices provide mappings from text phrases to point-level masks.
Figure 2: Main pipeline for Grounded 3D-LLM. The training process for Grounded 3D-LLM is divided into two key steps. Firstly, CLASP utilizes extensive scene-text annotations (at the phrase level) to pre-train a 3D point cloud encoder and a cross-modal interactor. The subsequent step involves multi-task instruction tuning, which interlaces referent tokens within the instructions and responses, thereby facilitating flexible 3D scene understanding tasks.
Step 2: Instruction tuning with scene referent tokens. During instruction tuning, the model learns to use referent token as the noun phrase to reference the scene, represented by a special token <ref> in the language. The language input and output formats are demonstrated in the following example:
[All Scene queries] USER: Describe the object <ref> in the scene.
ASSISTANT: <p> A black chair </p> <ref> with four legs.
In this format, [All scene queries] derive from projecting Q-Former’s query features into the language embedding space. For input, <ref> is filled with the soft query embeddings via instance-query mapping. For output, <ref> serves as a placeholder token, aligning its hidden embeddings $R_{text}$ with the corresponding scene entities (single or multi-instances) as calculated by $S_{text}$:
@@
During training, the contrastive loss for $S_{ref}$ is applied as $S_{text}$ in the pre-training stage. During inference, referent tokens are decoded with a small decoder (two MLP layers) that extracts mask embeddings, retrieves queries and obtains their associated masks. Through the above definition of <ref> in language inputs and outputs, it enables a closed-loop dialogue with grounding in the 3D scene. To enhance the readability of the referent tokens, the model is trained to generate relevant text phrases followed by the tokens, such as “<p> three nearby chairs </p> <ref>.” special token pairs <p> and </p> signify the start and end positions of the noun phrases [58, 81]. If no category is provided, we use the generic term “object” instead.
Task-specific instruction templates. The flexible design of referent tokens enables the unification of downstream 3D vision tasks into language forms. We transform existing 3D vision tasks into instruction formats, as depicted in Fig. 1. This format encompasses single and multi-object grounding [8, 83], object detection [16, 60], dense captioning [14], 3D QA [4], and descriptions of objects and local scenes. For each task, we utilize a range of task-specific instruction templates (about 10-20), tailored to the data annotation formats. For tasks requiring grounding or detection, we append “(with grounding)” to instructions to indicate the need for generating referents in language outputs. These templates can be found in the supplementary materials.
Discussion: Referent token as localization token. Prior work 3D-LLM [31] employs discrete location tokens, which constrain the granularity of localization accuracy. Moreover, it lacks point-level grounding capabilities and struggles with grounded visual chat tasks. Chat-3D v2 [33], pairing object identifiers and object features as text inputs, confronts high token costs and dependency on independent detectors. Related works in 2D LLM [42, 81] rely on downstream expert models [41] to improve localization. Differently, our model operates without a separate detector and object feature extractor, relying instead on CLASP that aligns semantics with scene queries. It enables phrase-to-region mapping and associated mask retrieval. Additionally, each referent token can be decoded into multiple instances, thereby reducing token costs.
3.3 Grounded Language Data
In this work, we define “grounded language” as language annotations that exhibit phrase-to-region correspondence, as detailed in Sec. 3.1. We adopt two approaches to generate grounded language data: (i) automatic generation of grounded scene captions, and (ii) transformation of existing language data into a grounded format. Compared to existing datasets [8, 83, 2], the automated method offers two advantages: First, the captioning employs all objects as anchors to ensure complete coverage of 3D scenes. Second, the generated caption has a high object density, averaging one instance per nine tokens, as shown in Table 2, ensuring the training efficiency of positive instances per text.
Figure 3: Grounded scene caption dataset generation pipeline. Matching colors are used to link phrases with their corresponding object masks.

### v2

@@ Abstract
Prior studies on 3D scene understanding have primarily developed specialized models for specific tasks or required task-specific fine-tuning. In this study, we propose Grounded 3D-LLM, which explores the potential of 3D large multi-modal models (3D LMMs) to consolidate various 3D vision tasks within a unified generative framework. The model uses scene referent tokens as special noun phrases to reference 3D scenes, enabling it to handle sequences that interleave 3D and textual data. Per-task instruction-following templates are employed to ensure natural and diversity in translating 3D vision tasks into language formats. To facilitate the use of referent tokens in subsequent language modeling, we provide a large-scale, automatically curated grounded scene-text dataset with over 1 million phrase-to-region correspondences and introduce Contrastive Language-Scene Pre-training (CLASP) to perform phrase-level scene-text alignment using this data. Our comprehensive evaluation covers open-ended tasks like dense captioning and 3D question answering, alongside close-ended tasks such as object detection and language grounding. Experiments across multiple 3D benchmarks reveal the leading performance and the broad applicability of Grounded 3D-LLM. Code and datasets are available at the project page.
Index Terms:
Large Multi-modal Model, 3D Scene Understanding
@@ 1 Introduction
However, the exploration of a unified model for 3D perception and reasoning remains scarce, leaving it as an open problem. In our study, by leveraging the generative modeling capability of large language models (LLM), we introduce Grounded 3D-LLM for integration of multiple 3D vision tasks in linguistic formats (Fig. 1).
Its model design is inspired by the natural text, ”Two brown chairs are near the white nightstand,” which contains two noun phrases: ”two brown chairs” (region) and ”the white nightstand” (singular object). We refer to their corresponding 3D objects in the scene as ”phraseable objects,” indicating that each can be described in a text phrase and that together they are interrelated within the contextual description. Without loss of generality, we aim to represent these phraseable objects using a special token, termed the ”referent token” <ref>. Each referent token supports decoding to identify the corresponding objects (singular or plural) in a 3D scene. In the above example, <ref> can refer to items like “the white nightstand” or “two brown chairs”, substituting the corresponding text tokens in the language. The use of referent tokens as manageable noun phrases in linguistic descriptions offers two advantages: (i) a referent token can serve as either an alternative to or a companion for text phrases without altering the contextual description, as depicted in Fig. 1. and (ii) it enables the localization of nouns or text phrases within the language. Thus, this design positions referents as a unified interface, facilitating 3D grounding capabilities across various 3D vision tasks.
To accomplish this, the language modeling of the referent involves two high-level steps: (i) A more granular, phrase-level contrastive pre-training of the vision-language encoder in a large-scale scene-text dataset rich in phrase-level correspondence, and (ii) referent-aware instruction-following fine-tuning, which associates referent tokens with their respective object embeddings.
@@
Figure 1: Unifying scene-level 3D vision tasks via task-specific prompts in Grounded 3D-LLM framework. The interleaved referent token <ref> in language modeling allows the identification of external scene referents through masked decoding in raw point clouds. Words within <p> and </p> denote the accompanying text phrases. Matching colors are shown in the figure to connect phrases with their respective object masks.
Specifically, unlike CLIP, which aligns text with the whole image, we first perform contrastive alignment between text primitives – noun phrases in the contextual description – with their corresponding visual embeddings from 3D scene encoder. This scene-level pre-training process is referred to as Contrastive LAnguage-Scene Pre-training (CLASP). However, existing paired scene-text data at the phrase level (referred to as grounded scene-text data in this paper) remain scarce. To address this, we develop an automated pipeline to synthesize grounded scene-text data using a densely annotated 3D detection dataset. The resulting dataset, called Grounded Scene Caption (G-SceneCap), comprises 107K scene captions with over 1+ million phrase-level correspondence. Compared to existing scene-text datasets, it provides richer phrase-level correspondence, averaging one object every nine tokens, as shown in Table II. Besides, existing human-annotated 3D vision-language datasets [5, 23] are also into grounded language forms.
After the phrase-level scene-text alignment, we conduct multi-task instruction fine-tuning for Grounded 3D-LLM to associate the referents with their respective phraseable objects. We convert various existing grounded scene-text datasets into instruction-following format, where referent tokens are interleaved as alternatives or companions to the text phrases (examples shown in Fig. 1). During training, these special referent tokens are supervised to decode into specified object embeddings, identifying their external scene referents. These task-specific instruction-following conversions encompass a range of existing 3D vision-language tasks – from single to multi-object grounding, instance segmentation (object detection), to 3D QA and dense captioning. We compare this model’s prompt support and task versatility with previous large multi-modal models in Table I. We believe this represents a significant step toward unifying various 3D vision tasks within generative language modeling.
@@
  • We propose Grounded 3D-LLM, which establishes correspondence between 3D scenes and natural language using referent tokens. This method facilitates scene referencing and effectively models various 3D vision-language problems within a unified language modeling framework.
  • We first provide an automated, curated grounded scene caption dataset with over 1 million phrase-level correspondences. Experiments show that CLASP, using this data in both supervised and zero-shot text settings, demonstrates broad generalization in phrase-level grounding.
  • Without requiring specialized models or task-specific fine-tuning, the single model Grounded 3D-LLM achieves top-tier performance in most downstream tasks among generative models, particularly in grounding problems.
TABLE I: Comparison of 3D multi-modal models. We refer to instance segmentation as Inst.Seg., object box detection as Obj.Det., single-object grounding as Grd., point-level grounding as Point-Grd., multi-object grounding as Multi-Obj Grd., question answering as QA., and dense captioning as Cap..
@@
Text Vision Inst.Seg. Obj.Det. Grd. Point-Grd. Multi-Obj Grd. QA. Cap.
@@ 2 Related Works
2D large multi-modal models. Vision-language models [27, 22] are proposed to bridge visual and textual models via image-level contrastive learning. Their direct application enables open-world understanding [28, 29]. GLIP and its subsequent works [30, 31, 32] have innovatively formulated object detection as grounding problems, leveraging detection or pseudo-labels to align semantics. With stronger language models, various powerful large multi-modal models are capable of handling tasks including image captioning [33, 34], VQA [35, 36, 37, 38], and object detection/segmentation [32, 39]. Besides, incorporating multi-modal tokens into language modeling has been validated, including visual understanding [34, 40, 41, 42], robot control [43], and interleaved text-image content generation [44, 45]. On the other hand, Chen et al. [46] demonstrates that existing 2D VLMs lack 3D spatial reasoning, primarily due to insufficient 3D spatial data.
@@
3D scene understanding. In 3D scene understanding, prior works [47, 48, 49] have advanced large-scale unsupervised learning for foundational models. Language is increasingly used to encapsulate user intentions due to its broad applicability, ranging from closed-set object detection [1, 2, 4, 50, 24] to open-set 3D detection [51, 52, 53, 54, 55, 56]. Typically, these 3D detectors lay the groundwork for subsequent scene-language tasks. Such tasks encompass 3D Visual Grounding [5, 57, 7, 23, 9, 58, 6, 59], where models pinpoint objects within a 3D scene based on language queries. They also include 3D Question Answering [10, 60, 61], which demands robust spatial perception and reasoning. Another task, 3D Dense Captioning [14, 62, 63], involves localizing and describing all scene objects, necessitating a comprehensive understanding of object locations and attributes. While initial efforts were often task-specific and lacked generalizability, some methodologies [17, 8] have integrated 3D visual grounding and dense captioning tasks, leveraging their complementary nature. Recent endeavors such as 3D-VisTA [25] and 3D-VLP [13] aim to establish a universal 3D visual-language framework by pre-aligning 3D scenes with language. In contrast, most approaches develop specialized models or require per-task training. They also rely on off-the-shelf 3D detectors.
@@
3D large multi-modal models. 3D object-level LMMs [64, 65, 66, 67] have effectively bridged the gap between 3D visuals and texts by leveraging extensive 3D object datasets [68]. This has advanced the development of 3D object-level LMM [69, 70, 21, 71, 72, 67, 73]. However, these models struggle with interpreting complex spatial interrelationships in 3D scenes. For scene-level LLMs, models such as Chat-3D [19], Scene-LLM [74], LL3DA [20], and LEO [75] enable scene-level dialogue, achieving notable effectiveness in tasks such as question answering and captioning. Chat-3D v2 [26] enables language grounding by using object identifiers for effective referencing. 3D-LLM [21] improves scene understanding by incorporating positional embeddings and location tokens. Our approach utilizes special referent tokens for flexible scene modeling and supports diverse tasks including object detection and grounding. Additionally, our model functions as a generalist model without requiring task-specific fine-tuning.
@@ 3 Method
This section consists of three parts. Sec. 3.1 examines vision-language alignment at three levels. Sec. 3.2 details the Grounded 3D-LLM, from vision-language pre-training to multi-task instruction fine-tuning with referent tokens. Finally, Sec.3.3 describes the construction of the large-scale grounded scene-text dataset for pre-training, while Sec.3.4 illustrates the conversion of existing scene-text data to instruction-following formats for LLM fine-tuning.
@@ 3.1 Preliminary: Vision-Language Correspondence
The significance of vision-language alignment in 2D foundation models, exemplified by the Contrastive Language-Image Pre-training (CLIP) [27], is well-established. We explore three levels of vision-language alignment:
  • Sentence-level Correspondence: CLIP models [27, 22] introduced contrastive learning between image and sentence-level text embeddings. However, the limited size ($\sim$100K-scale) of paired scene-text datasets[5, 57, 23] hinders broader application, and the alignment lacks object-level grounding.
  • Word-level Correspondence: Initially introduced by GLIPs [30], it reinterprets object detection labels as phrase grounding tasks. However, this reliance on existing detection labels does not adequately capture appearance and spatial relationships.
  • Phrase-level Correspondence: As referred to as phrase-to-region correspondence in prior work [76], this constraint requires linking text phrases to specific objects (singular or plural) within a contextual description including category, attributes and etc.. However, such 3D datasets remain scarce, especially in 3D scenes.
Figure 2: Main pipeline for Grounded 3D-LLM. The training process for Grounded 3D-LLM is divided into two key steps. Firstly, CLASP utilizes an extensive grounded scene-text dataset (at the phrase level) to pre-train a 3D point cloud encoder and a cross-modal interactor. The subsequent step involves multi-task instruction fine-tuning, which interlaces referent tokens within the instructions and responses, thereby facilitating flexible 3D scene understanding tasks.
Our motivation. To enable the natural referencing of scene objects in language modeling, we aim to model phrase-level correspondences, allowing phraseable objects to be represented as referent tokens. In other words, referent tokens can substitute for or accompany noun phrases without altering the language structure. Inspired by Llava [35], we first adopt contrastive pre-training for the vision-language encoder at the phrase level using grounded scene-text data. During instruction-following fine-tuning, the LLM associates the newly proposed referent token with these phraseable objects. For example, in the phrase ”two brown chairs near the table,” the average word embeddings of ”two brown chairs” align with the 3D visual embeddings of the corresponding chairs. By inputting these text-aligned object embeddings, the LLM can reference these phraseable objects using the special referent. This can be formatted as <p>two brown chairs</p> <ref> near the table or simply as <p>object</p> <ref>. The token pairs <p> and </p> indicate phrase boundaries [39, 42].
@@ 3.2 Grounded 3D-LLM
Based on the aforementioned motivation, the main pipeline of Grounded 3D-LLM is illustrated in Fig. 2. To reference phraseable objects in language modeling, a vision-language model is first pre-trained on large-scale grounded scene-text data to align text phrases with their corresponding 3D objects. Subsequently, a large language model (LLM) is fine-tuned using multi-modal instruction-following data, where referent tokens serve as interleaved “soft” text tokens representing the phraseable objects. Per-task instruction-following templates are employed to address the diverse range of 3D vision tasks within the unified language modeling framework.
Step 1: Contrastive language-scene pre-training (CLASP). In the same spirit as CLIP [27], CLASP is pre-trained to align 3D scene embeddings with textual descriptions using grounded scene-text data. Each paired grounded scene-text annotation contains phrase-to-region correspondences, linking noun phrases inside the text and their respective 3D object IDs (singular or plural). CLASP co-trains a 3D encoder, a text encoder, a set of object queries, and a cross-modal interactor (CMI), as shown on the left side of Fig. 2. Specifically, a sparse convolutional network [77] serves as the 3D multi-level feature encoder, extracting point-wise embeddings $P$ as a $K$-level feature pyramid, while BERT [78] extracts text queries $T$. For each pyramid level, a cross-modal interactor (CMI) is used to extract object embeddings conditioned on both point cloud embeddings and their respective text descriptions. Specifically, CMI interacts with $N$ learnable object queries $Q$ as proxies to connect an $M$-word text (extracted from BERT) with point clouds. Initially, these object queries interact with $k$-level point-wise embeddings via cross-attention. Then, as shown in Fig. 3, to incorporate the contextual text description, they interact with text embeddings through plain one-way cross-attention or bi-directional cross-attention [32].
Figure 3: Designs of Cross-Modal Interactor (CMI). Sub-figure (b) shows the bi-directional cross-attention design inspired by GroundingDINO [32], which is the default choice.
Figure 4: Grounded Scene Caption dataset generation pipeline. This pipeline leverages privileged knowledge in 3D, including dense object labels, multi-view images, and global spatial relationships, along with the capabilities of ChatGPT tools and spatial relationship programming tools, to establish accurate scene-level captions with grounding information.
As shown in Fig. 2 (left), CMI yields both object and text embeddings, followed by a small head that transforms them into the textual embedding space: $P_{text}$ for positive text embeddings and $Q_{text}$ for object embeddings, as well as the mask embedding space: $P_{mask}$ and $Q_{mask}$ for mask embeddings of both, respectively:
$\left[P_{mask},Q_{mask}\right],\left[Q_{text},P_{text}\right]=\text{CMI}(P,T,Q).$ (1)
@@
As depicted in Fig. 2, we compute the alignment scores $S_{mask}$ for per-point mask classification and $S_{text}$ for the textual embedding similarities between noun phrases and object queries:
$S_{mask}=P_{mask}Q_{mask}^{\top},\quad S_{text}=P_{text}Q_{text}^{\top},$ (2)
@@
where we ignore the temperature $\eta=0.1$. The per-point mask classification for $S_{mask}$ is computed to classify each point against the object queries using bipartite matching [79, 80]. Another contrastive loss for noun phrases and object queries, $S_{text}$, is supervised by the phrase-to-region correspondence. During inference, the mapping from text phrases to point-level masks can be computed using both similarity matrices, e.g., by employing the top-scored indices as the mapping.
Step 2: Instruction fine-tuning with referent tokens. During instruction-following fine-tuning, Grounded 3D-LLM models input and output in a unified multimodal language format, fine-tuning a LLM based on the pre-trained CLASP. Specifically, the LLM inputs scene encoding (all object query embeddings) from the frozen CLASP with a plain detection prompt [30]. It aims to learn to use the referent token <ref> to reference the scene, given a large amount of grounded instruction-following data.
Given a single grounded instruction-following data point, the text is interleaved with referent and each referent has its mapping to some phrasesable objects (referent correspondence). We first convert the instruction and response into the USER-ASSISTANT format before feeding them into the LLM. An example is shown below:
[All object embeddings] USER: Describe the
object <ref> in the scene.
ASSISTANT: <p> A black chair </p> <ref>
with four legs.
To enable closed-loop dialogue with the LLM using scene referents, we model referents as their corresponding objects in both language input and output. Specifically, for language input, the referents (visible to the model) are directly replaced with their respective object query embeddings, followed by a linear projection to match the LLM channel size. For language output, the referents serve as placeholders, with their hidden embeddings $R_{text}$ aligned with the corresponding object query embeddings. During fine-tuning, we compute the similarity matrix between the referent and all object query embeddings $Q_{text}$:
@@
The contrastive loss for $S_{ref}$ is supervised with the referent correspondence and applied with the same temperature $\eta$ as CLASP. During inference, the output referents are decoded with a decoder (two MLP layers) to extract hidden embeddings $R_{text}$, retrieve the respective object queries, and obtain their associated masks through prior CLASP.
Besides, we ensure linguistic naturalness by formatting referents with accompanying text phrases, such as “<p> three nearby chairs </p> <ref>.”. If no category is provided in scene-text annotations, we use the generic term “object” instead. For tasks requiring grounding or localization, we append “(with grounding)” to instructions to enable referencing in the language output.
Discussion: Referent token as localization token. Prior work, 3D-LLM [21], uses discrete location tokens, limiting localization accuracy and hindering grounded visual chat tasks. Chat-3D v2 [26] pairs object identifiers with features as text inputs but incurs high token costs and relies on off-the-shelf detectors and per-object feature extractors [67]. In contrast, our model jointly extracts per-object features through large-scale, fine-grained scene-level pre-training of CLASP. The design of the Referent token allows decoding into multiple instances, reducing token costs and increasing localization efficiency. Additionally, it is the first to support point-level mask decoding.
3.3 Generating Grounded Scene-Text Data for Pre-training
In this subsection, we describe two approaches to generate grounded scene-text data for pre-training: (i) automatic generation of grounded scene captions, and (ii) transformation of existing language data into grounded formats.
Figure 5: Grounded Scene Caption visualization. In each grounded text, a phrase-to-region correspondence is formatted as “[object_phrase object_IDs]”, where “object_phrase” refers to a noun phrase and “object_IDs” denotes the corresponding object IDs in the groud-truth annotation. The bottom line demonstrates explicit phrase-to-region correspondence example between the noun phrases (character positions) and the corresponding objects.
TABLE II: Grounded scene-text dataset comparison. “Sent.”, “Obj.”, “Multi.Obj.”, and “Corr.” refer to “Sentence”, “Object”, “Multi-objects”, and “Correspondence”. Tokens are calculated using BERT [78]. For sentences with single-object correspondence, each text is paired with one instance.
Dataset #3D Scan #Tokens #Text #Tokens/Text Correspondence
#Corr. Type Corr/Token
ScanRefer [5] 0.7K 1.17M 52K 22.6 52K Sent.-Obj. 4.4%
Sr3D [57] 1.5K 1.10M 84K 13.1 84K Sent.-Obj. 7.7%
Sr3D+ [57] 1.5K 1.48M 115K 12.9 115K Sent.-Obj. 7.7%
Nr3D [57] 0.7K 0.617M 42K 14.7 42K Sent.-Obj. 6.8%
Multi3DRefer [23] 0.7K 1.20M 62K 19.4 87K Sent.-Mult.Obj. 7.2%
ScanScribe [25] 3.0K 18.5M 278K 66.5 278K Sent.-Obj. 1.5%
Grounded Scene Caption 1.5K 9.19M 107K 85.9 1.02M Phrase-Multi.Obj. 11.1%